OpenStudy (anonymous):

if you add 1/3, 1/4, and 1/8 of a number you get 585... what is the number?

hero (hero):

So in other words \(\frac{n}{3} + \frac{n}{4} + \frac{n}{8} = 585\) or more simply, \(n(\frac{1}{3} + \frac{1}{4} + \frac{1}{8}) = 585\)

OpenStudy (valpey):

\[\frac{1}{3}x+\frac{1}{4}x+\frac{1}{8}x = 585\]\[x(\frac{1}{3}+\frac{1}{4}+\frac{1}{8}) = 585\]

OpenStudy (valpey):

3,4, and 8 have a LCM of 24.

OpenStudy (anonymous):

and then what do you do after...?

OpenStudy (anonymous):

n(8/24+6/24+3/24)=585

OpenStudy (anonymous):

so n(17/24)=585

OpenStudy (anonymous):

thank you so much "peanutman" that really helped a lot, i will give you a medel :)

OpenStudy (anonymous):

585 how many 17/24...

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

i'm doing my math packet for school and it's sometimes challenging so thanks :)

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

i'll need it!!!